AO Commissioning on 12 June 2011

Observer(s): D. Miller, JC. Guerra, K. Boutsia, E. Pinna, L Busoni, A. Puglisi, S. Esposito, J. Hill (LBTO)

Telescope Operator: S. Allanson (LBTO)

SW Support: N. Cushing (Tucson)

Telescope Support: M. Midkiff (LBTO)

Instrument Support: T. Shih (LBTO)

Summary

A beautiful photometric night with 0.5-0.6 arcsec seeing (2 sec in r' on the guider). Most of the night was spent testing new features of the AO software plus refining the AO parameters. After a few hours of that, there were some very nice high-Strehl images. These tests included using off-axis reference stars in binary pairs We used 1.5 hours at the end of the night to try a deep science exposure on John's Q2051+1950 field.

Details

02:30 twilight flats

03:02 iif_preset, ACT0341, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

03:06 iif_preset, BS9158, MODE=ACTIVE, GS=0, PA=-160

03:20 iif_preset, Pisces_13, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

03:26 iif_preset, Pisces_13, MODE=ACTIVE, GS=3, PA=90

03:31 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=90
  • irc_runao -failed because R=4mag which was too bright - replaced R=9.5

03:38 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=90
  • irc_runao

03:48 loop was closed successfully BUT primary went down and the loop opened - Steve added IssueTrack 3229

03:54 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=90
  • irc_runao
  • iif_offset, 0, -10, COORD='DETXY'
  • iif_offset, 1, -14, COORD='DETXY' - the offset failed before the 'Resume' command - the sec.AO ripped for unknown reasons

04:05 iif_preset, Pisces_13, MODE=ACE, GS=3,RS=0, PA=90
  • irc_runao - failed - BCU error

Repeat that multiple times because the shell keeps ripping.........

04:22 irc_runao - success finally!!!
  • iif_offset, 1, -14, COORD='DETXY'

04:27 The GCS/AGw is exactly in focus for PISCES filter #8, FeII.

04:50 we wanted to offset the bright star outside the field - before doing that it was advisable to apply the collimation look up table from the PSF primary GUI in order not to loose AO loop on a resume after an offset - we turned on guiding with GCS and caused the loop to open.

04:53 iif_preset, Pisces_13, MODE=ACE, GS=3,RS=0, PA=90 - preset failed - the star was on the edge

04:54 iif_preset, Pisces_13, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

04:56 iif_preset, Pisces_13, MODE=ACE, GS=3,RS=0, PA=90
  • irc_runao - success
  • DIMM - 0.9''
  • iif_offset, 1, -14, COORD='DETXY' - in order to put the star in the field (for some reason after the acquisition it ends up outside)

05:43 we need to do offsets in open loop - for this reason we need to collimaete the primary mirror - we tried this and it failed - we need to preset again - because the secAO ripped.

05:47 iif_preset, Pisces_13, MODE=ACE, GS=3,RS=0, PA=90
  • irc_runao
  • now we are trying open loop offsets
  • iif_offset, 5, 5, COORD='DETXY'
  • iif_offset, -5, -5, COORD='DETXY'
  • iif_offset, 20, 20, COORD='DETXY' - failed before resume (again) - we are changing target

05:57 iif_preset, Pisces_14, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

06:00 iif_preset, Pisces_14, MODE=ACE, GS=3, RS=0, PA=180
  • irc_runao - success
  • we take data for testing non common path abberation - slope null

Zernike Offload Tests

06:34 - secAO ripped - send the preset again : iif_preset, Pisces_14, MODE=ACE, GS=3, RS=0, PA=180
  • irc_runao
  • go on with non common path abberation tests

06:55 iif_preset, Pisces_14, MODE=ACE, GS=3, RS=0, PA=180
  • we change the PA because we are reaching the limit for the LN
  • iif_preset, Pisces_14, MODE=ACE, GS=3, RS=0, PA=90
  • irc_runao
  • in the same field we want to test the offload on the primary mirror (apply zernikes directly on the primary and see if the AO registers the same values) - the signs of astigmatism match - the signs of coma are reverse.

Time Sent to M1 Zern Sent Seen on AdSec Zern Seen
7:26:00 -100 z6 too small --
7:27:30 -500 z5 -500 z5
7:29:10 +400 z7 -400 z7
7:32:40 +500 z6 +500 z6
7:33:30 -400 z8 +300 z8
7:34:33 +300 z9 +400 z9
7:35:05 -300 z10 -500 z10
7:36:00 +500 z11 +500 z11

07:45 Steve modified the file /lbt/tcs/current/aos/configuration/DXoffloadmodes.dat so corrects in meters will be sent, as a test. PSF expects nanometer

07:53 iif_preset, Pisces_14, MODE=ACE, GS=3, RS=0, PA=90
  • irc_runao
  • AOS successfully sending Z5-Z11 corrections to PSF. Sending them in meters as a first test and see the numbers are reasonable and have the correct signs. Reasonable means that if they were converted to nanometers, they would all be less than 10 nm.

08:07 iif_preset, Pisces_14, MODE=ACE, GS=3, RS=0, PA=90
  • irc_runao
  • AOS is now sending nm of offloads from the shell, with a scale factor of 0.2. Most corrections of order 30 nm.

08:13 Apply 500 nm z6 and z11. See about 100 nm steps. Can not see this this on the AdSec GUI

08:16 Apply 5000 nm of Z6. See this on AdSec Position display. Send ~1000 nm offload for 3 iterations. Position of shell goes back to original. Offload modes works!!

This allows us to close Issue 2757.

On-sky Shell Flat

08:24 Collect data to create on-sky flat from 5 minutes of AO Closed Loop on a bright star. Elevation 71.

WFS Optical Gain

08:37 Test of the wavefront sensor optical gain. A disturbance is applied to the secondary with one mode and the gain optimization procedure starts

08:56 The rotator faulted - the loop opened but the mirror did NOT rip

09:00 Aborted a preset as the AO guys weren't quite ready to go.

09:06 iif_preset, Pisces_14, MODE=ACE, GS=4, RS=0, PA=0
  • irc_runao (failed - Reading shared buffer masterdiagnostic.R:OPTLOOPBUF)

09:15 iif_preset, Pisces_14, MODE=ACE, GS=4, RS=0, PA=0
  • irc_runao - DIMM 1.0''

Discovered that the high modes noise was coming from a board in the flower pot which is not used for observing. Turning it off made a big improvement in the modal noise. Improved images starting around 297.

John's Observations of Q2051+1950 (a.k.a. B2048+196)

IDL won't accept the object name with a + sign, so we rename it QSO. (later QSO2051)

   0    20 51 12.661    +19 50 06.35     0.0    0.0   2000.0  J2000     0.00    0.00    0.00   Target QSO
   1    20 51 11.8407   +19 49 48.703   -5.1   -2.5   2000.0  J2000   12.377    2.01   21.11 Reference Star

Observation design

1 hour total integration in Ks.
Divided into 2 position angles 50 degrees apart (avoid 60 /120 deg symmetry of secondary)
Each half hour divided into 5 dither positions.
Take 6 x 60 sec exposures with PISCES at each dither position.

Observation

09:33 iif_preset, QSO, MODE=ACQUIRE, TS=1, GS=0, RS=0, ROTMODE=IDLE - (J.Hill object)
  • using filter Ks
  • choose PA to keep PISCES upright

09:40 iif_preset,QSO, MODE=ACE, GS=11, RS=1, PA=-70 - didn't find the guide star in a convincing manner.

09:41 correcting pointing - iif_preset,WT10_272, MODE=ACQUIRE, GS=0, RS=0, ROTMODE=IDLE

09:xx iif_preset,QSO, MODE=ACE, GS=11, RS=1, PA=-70
  • irc_runao -sucess
  • iif_offset, -5, -3, COORD='DETXY' - offset failed! We need to do closed loop offsets!

09:53 iif_preset,QSO, MODE=ACE, GS=11, RS=1, PA=-70
  • iif_offset, -5, -3, COORD='DETXY' - Refused - because it found the loop paused (and considered it closed) - we send acquire_ref anyway
  • irc_runao (and it worked because the reference star was still in the CCD47 field)

AO is running at 625 Hz, with binning=2, and correcting 153 modes.

PISCES is taking 6 x 60 sec exposures in Ks at each dither position (plus a couple test exposures at the beginning).
These are done with Alfio's closed loop script to dither and take PISCES images.
Five dither positions (on a 3 arcsec radius dice5) 0,0 3,0, 0,3 -3,0 0,-3.
Reference star is not in the PISCES field at this first Position Angle.
Seeing (without adaptive optics) in R-band is 0.6 arcsec on guider. Sky is photometric, perhaps slightly smoky.

IMPORTANT: Position Angles written in the PISCES headers are wrong by 90 degrees. (Use double stars observed earlier tonight to sort out the actual postion angles.)
Also note that the RA and DEC values are exchanged in the headers of these images. JMH: Based on the observations of binaries PISCES_13 and PISCES_14 the True Astronomical Position Angles on PISCES are 90 degrees more negative than what the presets and headers say. A preset to PA=90 gives a true PA of 0.

Each 60sec exposure has a mean count level of 1730+-30 ADU.
For the first dither position, the QSO is at pixel 353,245.
We believe the AO reference star is off to the lower left, but didn't verify this.
The QSO has peak counts (including background) of 5430 ADU. SIMBAD say the magnitude is K=15.7, J=16.7, V=18.5.
The QSO is 4 pixels FWHM (0.018 arcsec/pixel TBC) in image 326, while a perfect Airy pattern should be 2.9 pixels FWHM.

First dither (0,0) starts with image 325.
First data set ends with image 354.

WARNING: Beware that the delay between the closed loop dither offset and the first exposure was shorter than expected. This means that the first image in each dither position could be offset/trailed. We don't see any obvious evidence of this in quick-look reductions by Juan Carlos. JMH: Yes, in data reduction, we see double images in images 349, 363 and 381, so the effect is there.

10:39 iif_preset,QSO, MODE=ACE, GS=11, RS=1, PA=-20 (Changing Position Angle by 50 deg)
  • irc_runao -sucess

Peculiar long integration (600 sec?) to start the next set. What causes this?? This is image 355 which appears to be a normal 60 sec exposure as viewed on ds9. So what was PISCES doing for 10 minutes? The sky/background counts in image 355 suggest that it was indeed a normal 60 sec exposure.

Image 356 is 10 sec to check the QSO position. It is the nominal positon of 225,486 pixels for the 0,0 dither.

357 is the first 60 sec exposure of the next set (PA=-20).
Second data set ends with image 386.
Twilight is starting to come up in the last ~10 exposures, but it doesn't affect the background levels in Ks.

back to AO tests on HR8799

11:30 iif_preset, Wref11, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

11:33 iif_preset, Wref11, MODE=ACE, GS=1, RS=0, PA=-90
  • irc_runao
 
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = Generic communication error (nobody replying or timeout)
        err = Resmsg[2]  = AcquireRef[54] failed - Generic communication error (nobody replying or timeout)

11:39 iif_preset, Wref11, MODE=ACE, GS=1, RS=0, PA=-90
  • irc_runao - success
  • maybe we are looking at the wrong star , we try pointing to the other star

11:49 iif_preset, Wref11, MODE=ACE, GS=0, RS=1, PA=-90

11:51 iif_preset, Wref11, MODE=TRACK, TS=1, GS=0, PA=-90
  • the background is getting too bright - end of science for tonight

11:55 iif_preset, WT10_304, MODE=ACQUIRE, GS=0, ROTMODE=IDLE - pointing for flats
  • filter J starting at image 413, exptime 10 sec
  • fillter H starting at image 417, exptine 8 sec
  • filter Ks staring at image 423 (8sec exptime) and the rest exptime 15sec
  • filter FeII starting at image 431, exptime 30sec and 15sec

12:20 - End of night

Sky flat summary

Image file    Filter.
Pisces.20110611.0009-.0012 Ks.
Pisces.20110611.0423-.0430 Ks
Pisces.20110611.0013-.0015 H.     
Pisces.20110611.0414-.0416  J
Pisces.20110611.0417-.0422 H
Pisces.20110611.0431-0440 FeII

-- KonstantinaBoutsia - 12 Jun 2011
Topic revision: r13 - 10 Feb 2012, MichaelWagner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback